Wages for desk clerks amount to $18,400/month.
Base charge for utilities is $3,340/month. An additional $1.80/night is incurred on average for each occupied room.
Fixed costs for taxes, insurance, advertising, etc. amount to $3,400/month.
The complimentary breakfast provided to guests costs on average $3.20 per occupied room.
Data regarding housekeeping costs for the last nine months are given below.
Cost ($) Rooms occupied
8,370 940
7,990 880
10,290 1,120
9,120 1,000
6,870 740
9,345 1,070
7,470 800
10,150 1,150
9,390 950
a) Using the hi-lo method, develop an equation for the behavior of housekeeping costs, in the form Y = a + bX
b) Develop an equation for total costs of operation, in the form Y = a + bX
c) Using your equation, forecast total operating cost if 1,250 rooms are occupied
d) Why might your answer to part c be unreliable?
e) What is the marginal cost of an occupied room? How might this information be of use to the owner?
Answer #1

Given data that Page | Cost(s) Rooms occupied 8,310 १५० 880 10,२१० 1,100 9,100 1,000 6,870 140 1,070 १,345 1,410 800 - 10,150page 2 a: 9=950+88 Vaspiable House Keeping cost peop stoom = $ (10,150 - 6870) / (11150-440) - $312801410 = $ 8 pesy goom Fixpage 3 di The Answest might be ungeliable since 19250 400ms is outside the gange psovided. e. The Marginal cost peg soom is $